OpenStudy (anonymous):

Shades R Us charges $20 per day to rent a lounge chair and $15 per day to rent an umbrella. Dan and Lisa paid a total of $245 to rent a lounge chair and an umbrella each during their vacation. Lisa rented the chair and umbrella for 1 day less than Dan. The following equation represents this situation, where x represents the number of days Dan rented the lounge chair and umbrella. 20x + 15x + 20(x − 1) + 15(x − 1) = 245 What is the total amount Dan paid to rent the lounge chair and umbrella during his vacation? a. $80 b. $105 c. $140 d. $175

OpenStudy (anonymous):

@iGreen @geekfromthefutur

OpenStudy (geekfromthefutur):

x(20+15+20+15) - (20+15) =245 70x - 35 =245 Add 35 to both sides and you'll get 70x -35 +35 = 245 +35 70x = 280 Then divide both sides by 70 and you'll get x =___?????? If you dont understand this et me know and i would do it slowly with you.

OpenStudy (geekfromthefutur):

Basic answer its D

OpenStudy (anonymous):

ok thanks that helped a lot

OpenStudy (geekfromthefutur):

Thats good let me know if you need help with any others

OpenStudy (igreen):

Geek got it :D